OpenStudy (anonymous):

The product of two consecutive negative integers is 288

OpenStudy (tfht):

-x(-x-1)=288

OpenStudy (tfht):

wait... i dont think that's possible

OpenStudy (anonymous):

x(x+1)=288, if x is the first integer. Then expand to get x^2+x=288. Rearrange the equation to get x^2+x+288=0, factor the left side and you will get 2 roots to the equation. You can reject the positive root since x must be negative. Therefore the negative root is x, and x+1 is the second integer, which is just the negative root +1.